Looking at the stoichiometry of the reaction (see the lab manual), how many moles KMnO4 must have been delivered by the buret to react with that much Fe(NH4)2(SO4)2 · 6H2O?
8H+ + 5Fe+2 + MnO4- --> Mn+2 + 5Fe+3 + 4H2O
(Remember that all Group I salts are soluble - including KMnO4. Therefore moles KMnO4 = moles MnO4-)

1.59×10−03 moles
5.92×10−04 moles
2.96×10−03 moles
5.03×10−04 moles

Considering the moles and volume of KMnO4 above, what must the true molarity of your KMnO4 solution be?

0.020 M
2.26×10−02 M
7.54×10−03 M
1.49×10−02 M
Answer #1

8H+ + 5Fe+2 + MnO4 - \rightarrow Mn+2 + 5Fe+3 + 4H2O

Mass of Fe(NH4)2(SO4)2 · 6H2O = 1.01 g

Molar mass of Fe(NH4)2(SO4)2 · 6H2O = 392.16 g/mol

Moles of Fe(NH4)2(SO4)2 · 6H2O = 1.01 / 392.16

= 0.0025

Moles of Fe+2 = 0.0025

From the equation:

5 moles Fe+2 reacts with 1 mole MnO4-

0.0025 moles Fe+2 will reacts with = 0.0025 / 5

= 0.0005 moles MnO4-

Therefore, moles of KMnO4 needed = 5.03x10-4 moles.

Volume of KMnO4 used = 25.68 - 0.15

= 25.53 mL = 0.025 L

Molarity of KMnO4 = 5.03x10-4 / 0.025

= 0.020 M